# ABACUS: An Electronic Structure Analysis Package for the AI Era

### v1

@@ III.5 Implicit Solvation Model
Electrochemical reactions, which refer to potential-driven processes at electrode/solvent interfaces, are becoming essential for developing green-energy technologies used in producing clean fuels. Understanding electrochemical interfaces’ structure and properties is crucial for designing and optimizing electrochemical systems. In this regard, atomic-scale computational simulations are pivotal in supporting the description of more complex electrochemical interfaces. However, theoretical modeling at the molecular level encounters significant challenges, such as addressing solvent layers, electrical double layers, and the variation of electron numbers at electrode/solvent surfaces, etc.
The implicit solvation model, together with the dipole correction and the compensating charge plate can achieve more realistic simulations of electrochemical reactions. These functions are integrated into the self-consistent field workflow of ABACUS, as illustrated in Fig. 9. After obtaining the initial electron density $\rho(\mathbf{r})$, several correction potential terms are evaluated, including the potential of the implicit solvent $V_{\mathrm{sol}}$, the compensating charge $V_{\mathrm{comp}}$, and the dipole correction $V_{\mathrm{dip}}$. These potential corrections are then applied to the Kohn-Sham equation, allowing computation of new electron density and the total energy.
@@
Figure 9: Implementation of the implicit solvation model in ABACUS. The self consistent loop of solving the Kohn-Sham equation starts with an initial electron density $\rho(\mathbf{r})$. Then, the external potential $V_{\mathrm{ext}}$, the Hartree potential $V_{\mathrm{H}}$, and the exchange-correlation potential $V_{\mathrm{XC}}$ are constructed. In the solvation model, the Hamiltonian consists of three additional terms (highlighted by red color), which are the implicit solvation potential $V_{\mathrm{sol}}$, the dipole correction term $V_{\mathrm{dip}}$, and the compensating charge term $V_{\mathrm{comp}}$. The corresponding energy corrections are also evaluated to yield the total energy $E[\rho]$. $E_{0}[\rho]$ denotes the system energy with no corrections.
@@
$S(\rho)=\frac{1}{2}\mathrm{erfc}{\left(\frac{\ln{\left(\frac{\rho}{\rho_{\mathrm{c}}}\right)}}{\sigma\sqrt{2}}\right)}.$ ) . (28)
The parameter $\rho_{c}$ is the charge density cutoff determining at what value of the electron density the dielectric cavity forms. $\sigma$ determines the width of the diffuse cavity. This assumption leads to a smooth variation of the relative permittivity from $\epsilon(\mathbf{r})$=1 of the solute to $\epsilon_{b}$ in the solvent.
@@
$\nabla\cdot\Big{[}\epsilon(\rho)\nabla\phi\Big{]}=-4\pi(N-\rho),$ (29)
@@
$V_{\mathrm{el}}=-\frac{d\epsilon(\rho(\mathbf{r}))}{d\rho(\mathbf{r})}\frac{{|\mathrm{\nabla\phi}|}^{2}}{8\pi},$ (30)
@@
$V_{\mathrm{cav}}=\tau\frac{d|\mathrm{\nabla S}|}{d\rho(\mathbf{r})},$ (31)
@@
where $\tau$ is the effective surface tension parameter. The two corrections (Eqs. 30 and  31) are collectively referred to as the implicit solvent-induced potential term $V_{\mathrm{sol}}$. The energy correction terms is
@@
We benchmark the accuracy of the implicit solvation implementation by calculating molecular solvation energies $E_{\mathrm{sol}}$, which is defined as the total energy differences between a solvated condition and a vacuum condition in ABACUS and comparing them against VASPsol-calculated values Mathew et al. (2019). We can see from Fig.10 that the solvation energy results from ABACUS agree well with those produced by the VASPsol package. The minor discrepancies between the solvation energies computed by the two methods may be attributed to the pseudopotential difference. We use projector-augmented-wave (PAW) potentials in VASPsol and the norm-conserving pseudopotentials in ABACUS.
Figure 10: Solvation energies (the total energy difference between a solvated condition and a vacuum condition) calculated by VASPsol and ABACUS for H2O, HCl, H2, and CH4.
ABACUS also support to add dipole correction and compensating charge when modeling surfaces. The periodic boundary conditions imposed on the electrostatic potential create an artificial electric field across a slab. By introducing an isolated slab-shaped density distribution $\rho(\mathbf{r})$ that is normal to the $z$-axis, a dipole correction Bengtsson (1999) is added to the bare ionic potential to compensate for the artificial dipole field within the context of periodic supercell calculations.
Modeling a constant-potential electrochemical surface reaction requires the adjustment of electron numbers in a simulation cell. Simultaneously, we must preserve the supercell’s neutrality under the periodic boundary conditions. Thus, a distribution of compensating charge needs to be implemented in the vacuum region of surface models when extra electrons are added to or extracted from the system. The compensating charge implemented in ABACUS follows the methodology developed by Brumme et al. in 2014 Brumme, Calandra, and Mauri (2014). We assume that the monopole with a total charge of $-n_{\mathrm{dop}}$ per unit cell is located at $z=z_{\mathrm{mono}}$. In this case, the effective potential $V_{\mathrm{mono}}(\mathbf{r})$ is added to the Hamiltonian, and the additional term $E_{\mathrm{mono}}$ is included in the total energy. Since $V_{\mathrm{mono}}(\mathbf{r})$ is independent of the electronic density, it is unnecessary to update in a self-consistent manner. We derive the correction on the ionic forces induced by the presence of the monopole, allowing for calculating the electronic structure and complete structural relaxation in the field-effect configuration.
Figure 11: Electrostatic energy of an electron along the direction (z) perpendicular to the Pt surface calculated by Quantum ESPRESSO and ABACUS Sun et al. (2024). The gray vertical dashed line represents the position of the compensating charge plate, and the blue dashed line corresponds to the dipole correction. A schematic structure plot of the testing interface model is aligned below the average electrostatic energy curves. Gray, red, and white spheres correspond to platinum, oxygen, and hydrogen atoms, respectively.
We output the electrostatic energy of an electron along the direction (z) perpendicular to the Pt surface for a testing system with an added electron number $N_{e}^{\mathrm{extra}}=0.2$ in the simulation cell. The testing system is a (3×3) Pt (111) surface slab composed of three atomic layers illustrated in Fig. 11. The modeled electrode surface contains 27 Pt atoms with one monolayer hydrogen coverage and six explicit water molecules. We place a compensating charge plate in the vacuum region above the water layer, and a dipole correction is also included in DFT calculations. We can see from Fig. 11 that the electrostatic energy results from ABACUS are in perfect agreement with those computed from Quantum ESPRESSO Giannozzi et al. (2020).
IV Methods in Plane Wave Basis
IV.1 Kohn-Sham Equation in Plane Wave Basis
In the system with periodic boundary conditions, the electronic wave functions can be expanded on the basis of plane waves, which are written as
where $\mathbf{G}$ and $\mathbf{k}$ represent the wave vectors of plane waves and sampling points in the Brillouin zone, respectively, $n$ denotes the band index, and $\{c_{n\mathbf{k}}(\mathbf{G})\}$ are the expansion coefficients of plane wave basis. The k-point sampling can be chosen such as the Monkhorst-Pack scheme Monkhorst and Pack (1976).
The Kohn-Sham equation described with plane wave basis takes the form of
where $\tilde{V}_{\mathrm{KS}}(\mathbf{G-G^{\prime}})$ is the plane wave represenation of the Kohn-Sham potential $\hat{V}_{\mathrm{KS}}$, $\varepsilon_{nk}$ is the eigenvalue of the Kohn-Sham equation, and $\delta_{\mathbf{GG^{\prime}}}$ is the Kronecker delta function.
The total energy can be calculated as
[…]
According to the Hellmann-Feynman theorem Hellmann (1935); Feynman (1939), the force acting on atom $I$ is defined as Eq. 23 and the stress is defined as Eq. 25. Consequently, when including a non-local part in the norm-conserving pseudopotential, the force on atom $I$ of type $\tau$ can be divided into three parts, namely
$\mathbf{F}_{I\tau}=\mathbf{F}_{I\tau}^{\mathrm{Ewald}}+\mathbf{F}_{I\tau}^{\mathrm{L}}+\mathbf{F}_{I\tau}^{\mathrm{NL}},$ (37)
where $\mathbf{F}_{I\tau}^{\mathrm{Ewald}}$ is the Ewald force. $\mathbf{F}_{I,\tau}^{\mathrm{L}}$ is the force contributed by the local part of the pseudopotential, and $\mathbf{F}_{I,\tau}^{\mathrm{NL}}$ is from the non-local part. The local potential term is given by
$\mathbf{F}_{I\tau}^{\mathrm{L}}=-i\Omega\sum_{\mathbf{G}}{\mathbf{G}e^{i\mathbf{G}\cdot\mathbf{R}_{I}}v_{\tau}^{\mathrm{L}}(\mathbf{G})\rho^{*}(\mathbf{G})},$ (38)
where $\Omega$ is the cell volume and $\rho(\mathbf{G})$ is the electron density in the plane wave basis, which is computed as
The non-local potential term is written as
$\mathbf{F}_{I,\tau}^{\mathrm{NL}}=$ $-2i\sum_{\mathbf{k},n}W(\mathbf{k})\sum_{\mathbf{G,G^{\prime}}}f(\epsilon_{n};\mu)c^{*}_{n\mathbf{k}}(\mathbf{G})c_{n\mathbf{k}}(\mathbf{G^{\prime}})$ (40)
×$\times\Big{[}e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}(\mathbf{G^{\prime}-G})v_{\tau}^{\mathrm{NL}}(\mathbf{k+G,k+G^{\prime}})\Big{]},$
where $v_{\tau}^{\mathrm{L}}$ is the local potential of atom type $\tau$, $W(\mathbf{k})$ represents the weight of the k point $\mathbf{(}k)$.
$\sigma_{\alpha\beta}=\sigma_{\alpha\beta}^{\mathrm{Ewald}}+\sigma_{\alpha\beta}^{\mathrm{Hartree}}+\sigma_{\alpha\beta}^{\mathrm{xc}}+\sigma_{\alpha\beta}^{\mathrm{T}}+\sigma_{\alpha\beta}^{\mathrm{L}}+\sigma_{\alpha\beta}^{\mathrm{NL}},$ (41)
where
$\sigma_{\alpha\beta}^{\mathrm{Hartree}}=-2\pi\sum_{\mathbf{G}\neq 0}\frac{|\rho(\mathbf{G})|^{2}}{G^{2}}\left[2\frac{\mathbf{G}_{\alpha}\mathbf{G}_{\beta}}{G^{2}}-\delta_{\alpha\beta}\right],$ (42)
$\sigma_{\alpha\beta}^{\mathrm{xc}}=-\delta_{\alpha\beta}\sum_{\mathbf{G}}[\epsilon_{\mathrm{xc}}(\mathbf{G})-v_{\mathrm{xc}}(\mathbf{G})]\rho^{*}(\mathbf{G}),$ (43)
$\sigma_{\alpha\beta}^{\mathrm{T}}=$ $\frac{2}{\Omega}\sum_{\mathbf{k},n}W(\mathbf{k})\sum_{\mathbf{G,G^{\prime}}}f(\epsilon_{n};\mu)$ (44)
$c^{*}_{n\mathbf{k}}(\mathbf{G})(\mathbf{k+G})_{\alpha}\delta(\mathbf{G,G^{\prime}})(\mathbf{k+G^{\prime}})_{\beta}c_{n\mathbf{k}}(\mathbf{G^{\prime}}),$
$\sigma_{\alpha\beta}^{\mathrm{L}}=\sum_{\mathbf{G},\tau}S_{\tau}(\mathbf{G})\Big{[}\frac{\partial v_{\tau}^{\mathrm{L}}(\mathbf{G})}{\partial(G^{2})}2\mathbf{G}_{\alpha}\mathbf{G}_{\beta}+v_{\tau}^{L}(\mathbf{G})\delta_{\alpha\beta}\Big{]}\rho^{*}(\mathbf{G}),$ (45)
and
$\sigma_{\alpha\beta}^{\mathrm{NL}}=$ $-\frac{2}{\Omega}\sum_{\mathbf{k}n}W(\mathbf{k})\sum_{\mathbf{GG^{\prime}}\tau}f(\epsilon_{n};\mu)c^{*}_{n\mathbf{k}}(\mathbf{G})c_{n\mathbf{k}}(\mathbf{G^{\prime}})$ (46)
×$\times S_{\tau}(\mathbf{G^{\prime}-G})\frac{\partial v_{\tau}^{\mathrm{NL}}(\mathbf{k+G,k+G^{\prime}})}{\partial\epsilon_{\alpha\beta}},$
where the structure factor is
$S_{\tau}(\mathbf{G})=\sum_{I\in\tau}e^{i\mathbf{G}\cdot\mathbf{R}_{I}}.$ (47)
IV.2 Iterative Diagonalization Methods
$\textbf{for}\ m=0\ \textbf{to}\ n\ \textbf{do}$ $\quad\Psi\leftarrow[\psi_{0}\ldots\psi_{m-1}]$ $\quad L\leftarrow\Psi^{T}\psi_{m}$ gemv $\quad\psi_{m}\leftarrow\psi_{m}-\Psi L$ $\quad HP\leftarrow H\psi_{m}$ FFT & Nonlocal $\quad\lambda\leftarrow\bra{\psi_{m}}H\ket{\psi_{m}}$ dot repeat $\qquad g\leftarrow K(HP-\lambda\psi_{m})$ vecdiv, dot, vecadd $\qquad\Psi\leftarrow[\psi_{0}\ldots\psi_{m}]$ $\qquad L\leftarrow\Psi^{T}g$ gemv $\qquad g\leftarrow g-\Psi L$ $\qquad d\leftarrow g+\gamma\ d_{\text{old}}$ dot, vecmul, vecadd $\qquad h\leftarrow Hd$ FFT & Nonlocal $\qquad\psi_{m}\leftarrow\psi_{m}\cos\theta+d\sin\theta$ dot, vecadd $\qquad HP\leftarrow HP\cos\theta+h\sin\theta$ vecadd $\qquad\lambda\leftarrow\bra{\psi_{m}}H\ket{\psi_{m}}$ dot $\quad\textbf{until}\ convergence$ end for
Figure 12: Preconditioned conjugate gradient algorithm in ABACUS. Here, $\psi_{i}$ represents the current approximation of the eigenvector being updated, and $\lambda$ denotes the corresponding eigenvalue, known as the Rayleigh quotient. To enhance computational efficiency, we employ a preconditioner $K$ to approximate the inverse of the Hamiltonian matrix $H$. Adjacent to each step of the algorithm are the numerical kernels.
With the plane wave basis set, the KS equation is routinely solved using iterative diagonalization methods, such as Conjugate Gradient (CG) and Davidson (Dav) methods, because the basis number is typically large. In general, diagonalization of given Hamiltonian is the most time-consuming part in most KS-DFT calculations, so efficient methods with high-performance-computing devices are required.
Heterogeneous computing devices, such as graphics processing units (GPUs) and deep computing units (DCUs), have significantly advanced the computational capabilities of modern scientific applications, especially in computational chemistry and materials science Jia et al. (2013); Maintz and Wetzstein (2018); Giannozzi et al. (2020). The ABACUS software harnesses these advancements to enhance the efficiency and scalability of its plane-wave basis set calculations.
As an example, the CG’s procedure is shown in detail in Fig. 12. The implementation fully leverages the heterogeneous framework of ABACUS, running the same code set across different devices through the abstraction layer. Operators like gemv and vecdiv are encapsulated under a unified interface for various platforms, including CPUs, GPUs, and DCUs. This allows the algorithm to enjoy cross-platform high performance without focusing on the underlying implementation details of these basic operations.
ABACUS utilizes a heterogeneous framework to execute diverse code efficiently. The framework consists of two primary components: Math Libraries and Kernel Operators. Math Libraries provide a unified interface for linear algebra calculations and fast Fourier transforms (FFT) across various accelerators. These libraries include CUDA-accelerated implementations like cuBLAS, cuSolver, and cuFFT; ROCm-accelerated implementations such as hipBLAS, hipSolver, and hipFFT; and CPU libraries like Intel Math Kernel Library. Kernel Operators serve as an abstraction layer that simplifies the execution of custom kernels and memory management across different heterogeneous frameworks, including NVIDIA CUDA, AMD ROCm, and Intel oneAPI. This abstraction enables ABACUS to write code independent of specific accelerators, eliminating redundant code for different processor types and enabling execution on CPUs. Consequently, this framework enhances code portability and fully leverages the strengths of various hardware architectures to improve computational performance. In summary, this framework allows us to write code once and execute it on multiple hardware platforms.
The allocation (explicit memory requests) and deallocation (release of memory) of GPU memory are time-consuming processes. Therefore, minimizing these operations in performance-critical sections of the code is crucial. A naive approach is extending the lifespan of variables or making them global, which can compromise the program’s architecture. Instead, ABACUS introduces another abstraction layer by using custom memory allocators optimized for specific use patterns. These allocators recycle large memory chunks across different objects, avoiding the costly system-level memory release. According to our benchmarks, 90% of the execution time and memory usage occur on the accelerators during the SCF loop, with minimal control flow on the Host CPU.
Additionally, our PW GPU code supports fully single-precision calculations during the SCF loop, maximizing the use of common hardware. Single-precision calculations significantly reduce memory usage and computational load, making them particularly advantageous for large-scale simulations. This feature not only accelerates the SCF loop but also ensures compatibility with a broader range of GPUs, including those with limited double-precision capabilities. ABACUS achieves higher efficiency and performance by optimizing resource utilization and facilitating more extensive and detailed scientific computations.
To evaluate the efficiency of heterogeneous acceleration in ABACUS, we performed SCF calculations using plane-wave basis sets on diverse hardware platforms for systems with varying numbers of atoms. We compared CPU and DCU hardware from the Bohrium cloud platform boh (2024) with roughly the same compute time cost for a relatively fair comparison. In Fig. 13, we choose the Davidson method since the Davidson method is generally faster than the Conjugate Gradient method. By analyzing the computation times for the different test cases, we discovered that the acceleration ratio achieved by the DCU computations increases as the number of atoms in the system rises. For instances with approximately 50 atoms, the acceleration ratio achieved on the DCU platform can be up to 5-fold, significantly reducing the computational cost.
@@ IV.2 Iterative Diagonalization Methods
Figure 13: For a collection of examples with different numbers of atoms, the Bar charts represent the calculation time per iteration by using the CG method on DCU (Red), the Davison method on CPU (Blue), and the CG method on CPU (Green) respectively. The line chart shows their ratio. One can download the examples from the link ope (2024b). We choose the hardware c32_m64 for CPU calculations and 4*DCU for DCU calculations.
IV.3 Stochastic DFT
IV.3.1 Formulas

### v3

@@ II.5 Implicit Solvation Model
Electrochemical reactions, which refer to potential-driven processes at electrode/solvent interfaces, play a pivotal role in advancing green-energy technologies for clean fuel production. A fundamental understanding of the atomic-scale structure and physicochemical properties of these interfaces is indispensable for the rational design and optimization of electrochemical systems. Atomic-scale computational simulations have emerged as a powerful tool for elucidating the complex nature of electrochemical interfaces. However, theoretical modeling at the molecular level faces substantial challenges, such as solvent layers, electrical double layers, and variations in electron numbers at electrode-solvent interfaces, among others.
To address these challenges, we implement a combined approach incorporating (i) an implicit solvation model to approximate solvent effects, (ii) a dipole correction scheme to mitigate spurious electrostatic interactions, and (iii) a compensating charge plate to maintain charge neutrality. As illustrated in Fig. 8, these methodological components are integrated into the SCF workflow of ABACUS, where the potential of the implicit solvent $V_{\mathrm{sol}}$, the compensating charge $V_{\mathrm{comp}}$, and the dipole correction $V_{\mathrm{dip}}$ is included in iterative electron density and total energy calculations. This integrated approach enables more realistic simulations of electrochemical systems while maintaining computational efficiency.
@@
Figure 8: Implementation of the implicit solvation model in ABACUS. The self-consistent loop of solving the KS equation starts with an initial charge density $\rho(\mathbf{r})$, followed by the evaluation of correction potentials, including the implicit solvent potential $V_{\mathrm{sol}}$, the compensating charge potential $V_{\mathrm{comp}}$, and the dipole correction potential $V_{\mathrm{dip}}$. These corrections are subsequently applied to the KS equations for iterative electron density and total energy calculations, where $E_{0}[\rho]$ denotes the energy without corrections.
@@
[…]
The parameter $\rho_{c}$ is the charge density cutoff, defining the electron density at which the dielectric cavity forms. The parameter $b$ determines the width of the diffuse cavity. This assumption leads to a smooth variation of the relative permittivity from $\epsilon(\mathbf{r})$=1 of the solute to $\epsilon_{b}$ in the solvent.
@@
$\nabla\cdot\Big{[}\epsilon(\rho)\nabla\phi(\mathbf{r})\Big{]}=-4\pi\Bigl{[}N(\mathbf{r})-\rho(\mathbf{r})\Bigr{]},$ (26)
@@
$V_{\mathrm{el}}=-\frac{d\epsilon(\rho)}{d\rho}\frac{{|\mathrm{\nabla\phi}|}^{2}}{8\pi},$ (27)
@@
$V_{\mathrm{cav}}=\tau\frac{d|\mathrm{\nabla S}|}{d\rho},$ (28)
@@
where $\tau$ is the effective surface tension parameter. The two corrections (Eqs. 27 and  28) are collectively referred to as the implicit solvent-induced potential term $V_{\mathrm{sol}}$, while the energy correction terms are
$E_{\mathrm{el}}=-\frac{1}{8\pi}\int{\epsilon\left(\rho\right){|\mathrm{\nabla\varphi}|}^{2}\mathrm{d}\mathbf{r}}$ (29)
$E_{\mathrm{cav}}=\tau\int{|\mathrm{\nabla S}|\mathrm{d}\mathbf{r}}.$ (30)
We benchmark the accuracy of the implicit solvation implementation by calculating molecular solvation energies $E_{\mathrm{sol}}$, which is defined as the total energy difference between a solvated condition and a vacuum condition in ABACUS and comparing them against VASPsol-calculated values Mathew et al. (2019). We can see from Fig. 9 that the solvation energies from ABACUS agree well with those produced by the VASPsol package. The minor discrepancies between the solvation energies computed by the two methods may be attributed to the pseudopotential difference. Note that we use projector-augmented-wave (PAW) potentials in VASPsol and the norm-conserving pseudopotentials in ABACUS.
Figure 9: Solvation energies $E_{sol}$ (the total energy difference between a solvated condition and a vacuum condition) calculated by VASPsol and ABACUS for H2O, HCl, H2, and CH4 systems.
ABACUS also supports adding dipole correction and compensating charge when modeling surfaces, which were utilized in a recent work Sun et al. (2025). The periodic boundary conditions imposed on the electrostatic potential create an artificial electric field across a slab. By introducing an isolated slab-shaped density distribution $\rho(\mathbf{r})$ that is normal to the $z$-axis, a dipole correction Bengtsson (1999) is added to the bare ionic potential to compensate for the artificial dipole field within the context of periodic supercell calculations.
Modeling a constant-potential electrochemical surface reaction requires the adjustment of electron numbers in a simulation cell. Simultaneously, we must preserve the supercell’s neutrality under the periodic boundary conditions. Thus, a distribution of compensating charge needs to be implemented in the vacuum region of surface models when extra electrons are added to or extracted from the system. The compensating charge implemented in ABACUS follows the methodology developed by Brumme et al. Brumme, Calandra, and Mauri (2014). We assume that the monopole with a total charge of $n_{\mathrm{dop}}$ per unit cell is located at $z_{\mathrm{mono}}$ along the $z$ axis. In this case, the effective potential $V_{\mathrm{mono}}(\mathbf{r})$ is added to the Hamiltonian, and the additional term $E_{\mathrm{mono}}$ is included in the total energy. Since $V_{\mathrm{mono}}(\mathbf{r})$ is independent of the electron density, it is unnecessary to update in a self-consistent manner. We also implemented the correction on the ionic forces induced by the presence of the monopole, allowing for calculating the electronic structure and complete structural relaxation in the field-effect configuration.
Figure 10: Electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface calculated by Quantum ESPRESSO and ABACUS Sun et al. (2025). The gray vertical dashed line represents the position of the compensating charge plate, while the blue dashed line corresponds to the dipole correction. A schematic structure plot of the testing interface model is aligned below the average electrostatic energy curves. Gray, red, and white spheres correspond to platinum, oxygen, and hydrogen atoms, respectively.
We output the electrostatic energy of an electron along the $z$ direction perpendicular to the Pt surface for a testing system with an added electron number $N_{e}^{\mathrm{extra}}=0.2$ in the simulation cell. The testing system is a (3×3) Pt (111) surface slab composed of three atomic layers, as illustrated in Fig. 10. The modeled electrode surface contains 27 Pt atoms, with a monolayer of hydrogen coverage and six explicit water molecules. We place a compensating charge plate in the vacuum region above the water layer, and a dipole correction is also included in DFT calculations. Fig. 10 shows that the electrostatic energy results from ABACUS are in perfect agreement with those computed from Quantum ESPRESSO Giannozzi et al. (2020).
III Methods in Plane Wave Basis
III.1 Kohn-Sham Equation in Plane Wave Basis
In the atomic system with periodic boundary conditions, the electronic wave functions can be expanded wit the plane waves, which take the form of
where $\mathbf{G}$ and $\mathbf{k}$ respectively represent the wave vectors of plane waves and sampling points in the Brillouin zone, $n$ denotes the band index, and $\{c_{n\mathbf{k}}(\mathbf{G})\}$ are the expansion coefficients of plane wave basis. The $k$-point sampling method is typically employed as the Monkhorst-Pack scheme Monkhorst and Pack (1976). Next, the Kohn-Sham equation described with the plane wave basis can be written as
where $\hat{V}_{\mathrm{KS}}(\mathbf{G-G^{\prime}})$ is the plane wave represenation of the Kohn-Sham potential $\hat{V}_{\mathrm{KS}}$, $\varepsilon_{nk}$ is the eigenvalue of the Kohn-Sham equation, and $\delta_{\mathbf{GG^{\prime}}}$ is the Kronecker delta function. In this regard, the total energy can be calculated as
$E_{\mathrm{tot}}=$ $\sum_{n\mathbf{k}}f(\varepsilon_{n\mathbf{k}})\varepsilon_{n\mathbf{k}}-\frac{1}{2}\int\int{\frac{\rho(\mathbf{r})\rho(\mathbf{r^{\prime}})}{|\mathbf{r}-\mathbf{r^{\prime}}|}\mathrm{d}\mathbf{r}\mathrm{d}\mathbf{r^{\prime}}}$ (33)
$-\int{V_{\mathrm{xc}}(\mathbf{r})\rho(\mathbf{r})\mathrm{d}\mathbf{r}}+E_{\mathrm{xc}}[\rho(\mathbf{r})]+E_{\rm II},$
According to the Hellmann-Feynman theorem Hellmann (1935); Feynman (1939), the force acting on atom $I$ is defined as Eq. 21 and the stress is defined as Eq. 23. Consequently, when including a non-local part in the norm-conserving pseudopotential, the force acting on atom $I$ of type $\tau$ can be divided into three parts, namely
$\mathbf{F}_{I}=\mathbf{F}_{I}^{\mathrm{Ewald}}+\mathbf{F}_{I}^{\mathrm{L}}+\mathbf{F}_{I}^{\mathrm{NL}},$ (34)
where $\mathbf{F}_{I}^{\mathrm{Ewald}}$ is the Ewald force. $\mathbf{F}_{I}^{\mathrm{L}}$ is the force contributed by the local part of the pseudopotential, and $\mathbf{F}_{I}^{\mathrm{NL}}$ is from the non-local part. The local potential term is given by
$\mathbf{F}_{I}^{\mathrm{L}}=-i\Omega\sum_{\mathbf{G}\neq\mathbf{0}}{\mathbf{G}e^{i\mathbf{G}\cdot\mathbf{R}_{I}}v_{\tau}^{\mathrm{L}}(\mathbf{G})\rho^{*}(\mathbf{G})},$ (35)
where $\Omega$ is the cell volume and $v_{\tau}^{\mathrm{L}}$ is the local potential of atom type $\tau$. Moreover, $\rho(\mathbf{G})$ is the electron density in the plane wave basis, which is computed as
Furthermore, the non-local potential term is written as
$\mathbf{F}_{I}^{\mathrm{NL}}=$ $-2i\sum_{n\mathbf{k}\mathbf{GG^{\prime}}}f(\epsilon_{n\mathbf{k}})c^{*}_{n\mathbf{k}}(\mathbf{G})c_{n\mathbf{k}}(\mathbf{G^{\prime}})$ (37)
$\times\Big{[}e^{i(\mathbf{G^{\prime}-G})\cdot\mathbf{R}_{I}}(\mathbf{G^{\prime}-G})v_{\tau}^{\mathrm{NL}}(\mathbf{k+G,k+G^{\prime}})\Big{]},$
where $v_{\tau}^{\mathrm{NL}}$ is the nonlocal potential of atom type $\tau$.
$\sigma_{\alpha\beta}=\sigma_{\alpha\beta}^{\mathrm{Ewald}}+\sigma_{\alpha\beta}^{\mathrm{Hartree}}+\sigma_{\alpha\beta}^{\mathrm{xc}}+\sigma_{\alpha\beta}^{\mathrm{T}}+\sigma_{\alpha\beta}^{\mathrm{L}}+\sigma_{\alpha\beta}^{\mathrm{NL}}.$ (38)
The formulas of the terms can be found in Ref. Liu and Chen, 2022b.
III.2 Iterative Diagonalization Methods
@@
Figure 11: Preconditioned conjugate gradient algorithm for $H\psi_{i}=\lambda\psi_{i}$, where $i=1,\cdots,n$ is band index. Here $\psi_{i}$ is the NN-dimensional wave function and $H$ stands for the $N\times N$ Hamiltonian. A preconditioner $K$ is applied to the residual $H\psi_{i}-\lambda\psi_{i}$ to obtain a preconditioned gradient $g_{i}$, which is then used to construct an improved search direction $d_{i}$ by Polak-Ribière method Polak (1971). Next, the new $\psi_{i}$ and $\lambda_{i}$ are obtained Teter, Payne, and Allan (1989).
@@
In general, diagonalization of the Hamiltonian matrix is the most computationally intensive step in most KS-DFT calculations. Therefore, the development of efficient diagonalization algorithms optimized for high-performance computing (HPC) architectures is essential. In particular, recent advances in heterogeneous computing platforms have substantially accelerated scientific computations, particularly in computational chemistry and materials science Jia et al. (2013); Maintz and Wetzstein (2018); Giannozzi et al. (2020). In some cases, the plane-wave basis set is capable of dealing with systems consisting of thousands of atoms or more Gygi et al. (2006); Watanabe et al. (2017); Lucignano et al. (2019); Gonze et al. (2020); Feng et al. (2024).
ABACUS employs a unified framework that operates seamlessly across diverse hardware architectures. Key linear algebra operations, such as generalized matrix-vector multiplication (gemv) and vector division (vecdiv), are encapsulated within a platform-agnostic interface, supporting execution on platforms such as CPU, GPU, and DCU. Consequently, the algorithm achieves high computational efficiency across multiple platforms. To evaluate the efficiency of heterogeneous acceleration in ABACUS, we conducted SCF calculations using plane-wave basis sets on the Bohrium cloud platform boh (2024), comparing CPU and DCU performance at equivalent computational costs. The Davidson method was selected for these benchmarks due to its demonstrated superiority in convergence speed compared to conjugate gradient approaches. Our analysis reveals a clear scaling relationship between system size and acceleration performance, with DCU computations achieving up to 5-fold speedup for systems containing approximately 50 atoms, as illustrated in Fig. 12. This performance improvement grows increasingly pronounced in larger systems, underscoring the effectiveness of our heterogeneous computing framework.
@@
Figure 12: Computation time per SCF step for a collection of examples (with different numbers of electrons) by using the conjugate gradient (CG) method on DCU (Red), the Davidson (DAV) method on CPU (Blue), and the CG method on CPU (Green), respectively. The line chart shows their ratio. The examples are available from the link ope (2024b). The hardware is c32_m64 for CPU calculations and 4*DCU for DCU calculations.
III.3 Stochastic DFT
III.3.1 Formulas